OpenStudy (anonymous):

solve 8x - (3x + 7) = 13 the solution is x = --?

OpenStudy (asadkarim7):

8x - 3x-7 = 13 5x = 13 + 7 20 / 5 = x x= 4

OpenStudy (saifoo.khan):

4

myininaya (myininaya):

i agree! :)

OpenStudy (anonymous):

i agree too :)

OpenStudy (anonymous):

8x-3x+7=13 5x=7+13 5x=20 x=20/5 =4